# Evaluating Impacts of Traffic Regulations in Complex Mobility Systems Using Scenario-Based Simulations

### v2

@@ 4.2 Inflow modifications
The absolute number of anticipated redistributed vehicles at each time interval is then computed as the fraction $F_{ar}(t)$ multiplied by the total anticipating $TN_{a}$. Similarly, the number of postponed redistributed vehicles is assessed by multiplying $F_{pr}(t)$ by $TN_{p}$.
@@
Given the proportion of vehicles implementing each behavioral strategy, it is possible to calculate the number of vehicles entering the policy-affected area under the modified scenario. Equation 14 presents the formulation for computing the inflow, which distinguishes between two cases depending on policy application time. Within the period of enactment of the policy, with $t\in[t_{s},t_{e}]$, the inflow is the sum of exempted vehicles and rigid vehicles. Outside the hours of the policy, with $t\notin[t_{s},t_{e}]$, the inflow is modified from a portion of vehicles that anticipate and postpone the starting time and sum to the original vehicles.
@@ 4.3 Traffic estimation model
In addition to the entrances, the policy affects the amount of traffic in the area. Our methodology does not attempt to model traffic variations directly; rather, its purpose is to capture the relationships between inputs and traffic, thereby enabling the estimation of modified traffic conditions given modified inflows.
@@
The total traffic load is generated by two sources: trips entering from outside the area (i.e., inflow) and trips starting within the area itself (i.e., starting trips). At any time interval $t$, one component of the traffic load consists of newly initiated trips, both inflow and starting trips, that begin during that interval. However, traffic does not reset at each time interval but evolves progressively as trips progress through the network. Many trips that are active at time t1t-1 remain active at time $t$, creating a temporal dependency between consecutive observations. This persistence means that the current traffic state is not independent from the previous one, but rather builds upon it. Consequently, the remaining traffic component is given by a fraction of the previous traffic load that persists into the current time step. Quantifying this persistence through a retention probability allows us to relate the evolution of traffic load to the pattern of trip departures over time.
@@
Given inflow and starting, we model traffic as a discrete-time system with observation intervals of $\nu$-minutes. At time $t=0$, the initial traffic in the observation area is $T(0)=I(0)+S(0)$, where $I(0)$ represents vehicles entering from outside, i.e., the inflow, and $S(0)$ represents vehicles starting their journey within the area. Then, for $t\in\mathbb{N}_{>0}$, traffic evolves according to the following Equation 15:
$T(t)=I(t)+S(t)+\alpha\cdot T(t-1)$ (15)
@@
where $\alpha$ is the retention probability, representing the fraction of vehicles circulating at time t1t-1 that will remain in the area at time $t$.
@@
Our approach models the retention probability, $\alpha$, based on assumptions about the dwell time in the area. Each trip is characterized by the duration for which it remains active within the area, i.e., the dwell time. We assume dwell times follow a geometric distribution with constant mean (Grimmett and Stirzaker, 2001). This choice is motivated by the memoryless property of the geometric distribution, which corresponds to assuming a constant per-step departure probability, independent of the elapsed dwell time. It also assumes that dwell times are independent of traffic conditions and that probabilities remain stable over time, regardless of congestion levels. Within this discrete-time framework, the retention probability $\alpha$ represents the probability that a trip does not end in a given time step and can be directly related to the mean dwell time.
@@
The objective is to compute the traffic vector satisfying Equation 15. Although this system admits a direct solution, we adopt an iterative approach as outlined in Algorithm 1. The direct and iterative approaches are mathematically equivalent and yield identical solutions; however, the iterative formulation empirically assesses superior computational performance, particularly in terms of execution time, and is hence chosen for the implementation. . The starting point is an initial guess for the traffic values, $T(t)^{0}=I(t)+S(t)$ for all $t$. Then, at each iteration step $k$, the algorithm recomputes the value of the retention probability $\alpha^{k}$ and updates the traffic estimate $T^{k+1}$, as shown in Equation 16.
$\begin{cases}T(0)^{k+1}=T^{0}(0)\\ T(t)^{k+1}=I(t)+S(t)+\alpha^{k}\cdot T(t-1)^{k}\qquad\forall t>0\\ \end{cases}\quad\forall k>0$ (16)
The estimation is repeated until the traffic vector $T^{*}$ satisfies the initial equation 15 given a certain tolerance $\gamma$. Specifically, convergence occurs when $\gamma$ is not overcome by the maximum difference of traffic in subsequent iterations, $\max_{t}|T^{(k)}(t)-T^{(k-1)}(t)|\leq\gamma$. A maximum number of iterations is also added to ensure a finite execution time.
@@
- Inflow and starting vectors $I,S\in\mathbb{R}_{\geq 0}^{n}$
@@
$n\leftarrow\text{length}(I(t))$ ;
/* Number of time intervals */
$T^{(0)}(t)\leftarrow I(t)+S(t)\quad\forall t\in\{1,\ldots,n\}$ ;
@@
$T^{(1)}(t)\leftarrow T^{(0)}(t)\quad\forall t\in\{1,\ldots,n\}$ ;
/* Copy initial state for iteration */
$k\leftarrow 1$ ;
/* Iteration counter */
while $k\leq K$ and $\max_{t}|T^{(k)}(t)-T^{(k-1)}(t)|>\gamma$ do
 $T^{(k+1)}(t)\leftarrow T^{(0)}(t)+\alpha^{(k)}\cdot T^{(k)}(t-1)\quad\forall t\in\{2,\ldots,n\}$ ;
@@
Algorithm 1 Traffic Estimation Algorithm
Having described the traffic model, we now address the computation of its input data, which are required to estimate the traffic vectors in both the reference and policy scenarios. First, in the reference scenario, $T$ is computed using the inflow and starting vectors $I$ and $S$. The inflows $I$ and the starting trips $S$ are extracted from observed data representing the current situation. Then, in the policy scenario, the modified traffic $T_{m}$ is estimated starting from the corresponding modified vectors $I_{m}$ and $S_{m}$, where $I_{m}$ is adjusted according to the procedures outlined in the previous subsection. For the starting vector, the same reasoning applied to the inflow can be adopted. All vehicles that start within the area of interest, excluding exempted vehicles, are required to pay a ticket fee and are therefore directly affected by the policy. As a result, they may decide whether and how to adjust their mobility habits with rigidity, time-shifting, mode-shifting, or cancellation. The modified starting vector was computed following the same procedure used for the inflow case, replacing the inflow vector $I$ with the starting vector $S$.
@@ 4.4 Emission estimation model
For the calculation of emissions in the area, it is assumed that each vehicle in circulation produces a value of average emissions per kilometer depending on its characteristics. This work uses the classification of vehicles in Euro emission standards (Selleri et al., 2021), widely adopted in European environmental and transport studies, to standardize the assessment of vehicle emissions and ensure comparability across different regions and datasets. Vehicles are classified into eight Euro classes, from Euro-0 to Euro-6, based on their characteristics such as vehicle type, fuel type, date of the first registration, and maximum emissions per kilometer. Accordingly, we define $\bar{E}_{l,km}$ as the representative average emissions per kilometer corresponding to the $l$-th Euro-class, with values obtained from the literature. Moreover, let $P_{l}$ represent the percentage of vehicles for each of the Euro-class $l$. This proportion reflects the actual composition of the fleet in the current base scenario, derived from the elaboration of real-world data.
The distribution of vehicles varies before and after the implementation of the policy, as the policy imposes different costs on different Euro-classes, which in turn affects the likelihood of vehicles being rigid. Consequently, during the period of application of the policy, the Euro-level distribution of the fleet must be recalculated considering the adjusted proportions of rigid vehicles, while outside this period it must be based on the original distribution. As a first step, the rigidities for each Euro-class must be calculated. Equation 17 represents the allocation of the total rigidity fraction, $F_{\text{rigidity}}$, among different Euro-levels $l$. In this way, the formula distributes the total rigidity among the partitions proportionally to both the size of the partition and the likelihood of rigidity within it for each time interval $t\in[t_{s},t_{e}]$.
$F_{rididity,l}(t)=F_{rigidity}(t)\cdot\frac{\mathbb{P}[\text{rigidity for Euro-level }l]}{\sum_{j}P_{j}\cdot\mathbb{P}[\text{rigidity for Euro-level }j]}$ (17)
Then, it is possible to compute the Euro-level distribution under the policy enactment. During the period $t_{s},t_{e}]$, the new proportion $P_{l,m}(t)$ accounts for two contributions. First, it considers the fraction of exempt vehicles $P_{l}\cdot F_{e}$, which represents the part of the Euro-class $l$ that is not subject to the policy and remains unchanged. Second, it considers the fraction of rigid vehicles, $P_{l}\cdot\frac{F_{rigidity,l}(t)}{F_{e}+F_{rigidity}(t)}$, which represents the proportional contribution of the rigid share of class $l$ compared to the total of rigid vehicles and exempt vehicles. Equation 18 formalizes this reasoning.
@@ 5.1 Technical development
The frontend is designed to ensure accessibility and usability for domain experts, including those without technical backgrounds. To this end, we develop an interactive dashboard that allows near real-time simulation of what-if scenarios under varying conditions. Users can directly modify key control parameters, described in Section 3 and 4, including policy fees per vehicle type, exempted vehicles, temporal windows of the policy application, behavioral parameters, and public transport settings. On the same page, simulation results are presented clearly and intuitively (see the next Section 5.3 for a wider discussion on qualitative and quantitative results). Global indicators summarize daily information for both the “as-is” baseline and “what-if” modified scenarios. They refer to indicators such as the total vehicle inflow per day, the maximum traffic within the day, total collected fees, and the estimated emission modification. For insight at a microscopic scale, interactive graphs are included to display the daily evolution of key quantities of interest in the current and modified scenarios. By changing the parameters, users can evaluate multiple scenarios corresponding to different combinations of parameters. The dashboard not only enables comparison between the baseline and modified scenarios, but also across different “what-if” scenarios (e.g., scenario 1 vs. scenario 2), supporting the assessment of alternative policy options.
@@ 5.2 Application: the case of Bologna, Italy
Figure 3: Left: the city of Bologna is highlighted in yellow, while the selected area, subject to the policy, is colored in orange. Center: population distribution in Bologna; the boundary of the selected area is highlighted in black. Right: road network and public transport stops in Bologna; the boundary of the selected area is highlighted in black
@@
The proposed model is tested on a real-world case study in the city of Bologna in Italy, a major urban center with a large resident population and a complex multimodal mobility system. The city is already characterized by a limited traffic zone in the historical center. For experimental purposes, we simulate an additional circulation restriction applied within a broader zone: its northern boundary is delineated by the ring road and the A1-A14 motorway, while the southern boundary follows the city’s morphological constraints, namely the hilly area. Figure 3 represents the selected area. This configuration serves as an illustrative example of how the model can be applied to large-scale area-based policies.
@@
The selected area exhibits both strong centrality and high urban attractiveness. It includes more than $321,172$ residents, 8282% of Bologna’s $392,044$ urban population [^1] Data extracted from the “Resident Population by Age, Sex, Citizenship, Neighborhood, and Statistical Area” dataset from the Comune di Bologna Open Data portal, version updated 31/12/2024; accessed 12/12/2025. URL: https://opendata.comune.bologna.it and 3131% of the $1,051,536$ residents in the wider metropolitan province [^2] Data extracted from the “Resident Population” ISTAT dataset; accessed on 12/12/2025. URL: https://demo.istat.it. It concentrates substantial private motorized mobility flows, with $269,226$ daily trips entering the area from outside and $329,807$ internal trips [^3] Origin-Destination data are extracted and relaborated from PUMS 2019, provided by Comune di Bologna. The road network of the selected area covers approximately $708$ km of drivable roads, of which $287$ km correspond to primary corridors (i.e., with tags trunk, primary, secondary, and tertiary roads according to OpenStreetMap) [^4] Road network data extracted from OpenStreetMap; accessed 8/06/2025. URL: https://www.openstreetmap.org. Public transport is also widely available in the area. In particular, for weekdays, the bus routes cover $249$ km of the drivable road network with $177$ bus lines, operating on a total of $786$ bus stops with a median frequency of $14$ trips per line per day (range: $1$ to more than 100100) [^5] Bus service data extracted from the Open Data portal of Trasporto Passeggeri Emilia‑Romagna (TPER); accessed on 25/05/2025. URL: https://solweb.tper.it/web/tools/open-data/open-data.aspx. The area also includes Bologna Centrale, a major national and regional railway hub. The station serves high-speed and intercity trains to major Italian cities (e.g., Florence, Milan, Venice, Rome) as well as international connections, including routes to Munich, and regional and suburban lines connecting surrounding towns and the Modena-Bologna-Ferrara corridor.
@@
Bologna is further selected due to the advanced mobility data infrastructure maintained by the municipality. High-resolution data are collected from multiple sources, including traffic sensors, cameras, and vehicle registration data from the national motorization authority. These datasets are already used for real-time traffic monitoring, allowing us to estimate the current “as-is” traffic situation. In particular, the integration of multi-source data with statistical and data science preprocessing techniques enables the estimation of inflows into the restricted area as well as trips originating within it, with a temporal discretization of $\nu=5$ minutes. Furthermore, we have estimates of the daily percentage of vehicles by Euro emission class currently circulating in the area. From Euro-0 to Euro-6 vehicles, respectively, the proportions are $0.059$, $0.012$, $0.034$, $0.054$, $0.198$, $0.176$, and $0.467$. Similarly, knowing the vehicular fleet composition in terms of segment, fuel type, and technology, we estimate the pollutant emissions per kilometer per vehicle typology in the area. We focus on the nitrogen oxides (NOx), a group of highly reactive gases emitted by vehicles that significantly contribute to air pollution. For each Euro-level class, the estimated emissions in grams per kilometer are: $0.2105847$, $0.217457$, $0.2401457$, $0.247239$, $0.135555$, $0.099559$, and $0.068246$.

### v3

@@ 4.2 Inflow modifications
The absolute number of anticipated redistributed vehicles at each time interval is then computed as the fraction $F_{ar}(t)$ multiplied by the total anticipating $TN_{ant}$. Similarly, the number of postponed redistributed vehicles is assessed by multiplying $F_{pr}(t)$ by $TN_{post}$.
@@
Given the proportion of vehicles assigned to each behavioral strategy, it is possible to calculate the number of vehicles entering the policy-affected area under the modified scenario. Equation 14 presents the formulation for computing the inflow, which distinguishes between two cases depending on policy application time. Within the period of enactment of the policy, with $t\in[t_{s},t_{e}]$, the inflow is the sum of exempted vehicles and rigid vehicles. Outside the hours of the policy, with $t\notin[t_{s},t_{e}]$, the inflow is modified from a portion of vehicles that anticipate and postpone the starting time and sum to the original vehicles.
@@ 4.3 Traffic estimation model
In addition to the entrances, the policy affects the amount of traffic in the area. Our methodology does not attempt to model traffic variations directly; rather, its purpose is to capture the dependence of traffic on selected quantities, specifically inflows and startings, allowing the estimation of modified traffic conditions when these quantities change.
@@
The total traffic load is generated by two sources: trips entering from outside the area (i.e., inflows) and trips starting within the area itself (i.e., starting trips). At any time interval $t$, one component of the traffic load consists of newly initiated trips, both inflow and starting trips, that begin during that interval. However, traffic does not reset at each time interval but evolves progressively as trips progress through the network. Many trips that are active at time t1t-1 remain active at time $t$, creating a temporal dependency between consecutive observations. This persistence means that the current traffic state is not independent from the previous one, but rather builds upon it. Consequently, the remaining traffic component is given by a fraction of the previous traffic load that persists into the current time step. Quantifying this persistence through a retention probability allows us to relate the evolution of traffic load to the pattern of trip departures over time.
Another key feature of traffic is its daily periodicity with characteristic recurrent patterns of morning and afternoon peaks and night-time lows. To formalize this periodicity, we define the set of observation times $\mathcal{T}=\{t_{0},t_{1},\dots,t_{n-1}\}$ as a discrete time space with equivalence modulo of 24 hours. This means that each integer time index $s\in\mathbb{N}$ can be associated with a representative $t\in\mathcal{T}$ such that $t=s\bmod n$, where $n$ is the number of observation intervals in a day. This construction ensures that operations like t1t-1 are always well-defined: for example, the interval immediately before the first observation after midnight corresponds naturally to the last interval of the previous day. In other words, the day is treated as a looped discrete-time system, simplifying the treatment of temporal dependencies at the boundaries of the day.
@@
Given the inflows and starting trips, we model traffic as a looped discrete-time system with observation intervals of $\nu$ (e.g., 5 minutes). At any time $t\in\mathcal{T}$, $I(t)$ represents vehicles entering from outside (i.e., the inflow) and $S(t)$ represents vehicles starting their journey within the area. Then, traffic evolves according to the following Equation 15:
$T(t)=I(t)+S(t)+\alpha\cdot T(t-1)\quad\forall t\in\mathcal{T},$ (15)
@@
where $\alpha$ is the retention probability, representing the fraction of vehicles circulating at time t1t-1 that will remain in the area at time $t$. This formulation combines both the incremental contribution of new trips and the persistence of existing traffic, while the looped time space $\mathcal{T}$ guarantees consistency across the limits of the day.
@@
Our approach models the retention probability, $\alpha$, based on assumptions about the dwell time in the area, i.e., the duration for which a vehicle remains active within the area. We make two key assumptions about vehicle dwell times. First, the probability that a vehicle departs at a given moment does not depend on how long it has already been waiting. Second, dwell times are assumed to be independent of traffic conditions, so the probability of departure is stable regardless of congestion. Given this modelling choices, we consider the dwell times to be geometrically distributed with constant mean (Grimmett and Stirzaker, 2001). Within this framework, the retention probability $\alpha$ represents the probability that a trip does not end in a given time step and can be directly related to the mean dwell time.
@@
The objective is to compute the traffic vector satisfying Equation 15. We adopt an iterative approach as outlined in Algorithm 1. The starting point is an initial guess for the traffic values, $T(t)^{0}=I(t)+S(t)$ for all $t\in\mathcal{T}$. Then, at each iteration step $k$, the algorithm recomputes updates the traffic estimate $T^{k+1}$, as shown in Equation 16:
$T(t)^{k+1}=I(t)+S(t)+\alpha\cdot T(t-1)^{k}\quad\forall t\in\mathcal{T}\quad\forall k>0.$ (16)
The estimation is repeated until the traffic vector $T$ satisfies the initial Equation 15 given a certain tolerance $\gamma$. Convergence is ensured by the fact that $\alpha<1$.
@@
- Set of observation interval $\mathcal{T}=\{t_{0},\dots,t_{n-1}\}$ - Inflow and starting vectors $I,S\in\mathbb{R}_{\geq 0}^{n}$
@@
$T^{(0)}(t)\leftarrow I(t)+S(t)\quad\forall t\in\mathcal{T}$ ;
@@
while $\max_{t}|T^{(k)}(t)-T^{(k-1)}(t)|>\gamma$ do
 $T^{(k+1)}(t)\leftarrow I(t)+S(t)+\alpha\cdot T^{(k)}(t-1)\quad\forall t\in\mathcal{T}$ ;
@@
Algorithm 1 Traffic Estimation Algorithm.
Having described the traffic model, we now address the computation of its input data, which are required to estimate the traffic vectors in both the reference and policy scenarios. First, in the reference scenario, $T$ is computed using the inflow and starting vectors $I$ and $S$. The inflows $I$ and the starting trips $S$ are extracted from observed data representing the current situation. Then, in the policy scenario, the modified traffic $T_{m}$ is estimated starting from the corresponding modified vectors $I_{m}$ and $S_{m}$, where $I_{m}$ is adjusted according to the procedures outlined in the previous Subsection 4.2. For the starting vector, the same reasoning applied to the inflow can be adopted. All non-exempt vehicles starting within the area of interest are subject to the ticket fee imposed by the policy, and their behavioral response is represented through the four adjustment strategies of rigidity, time-shifting, mode-shifting, or trip cancellation. The modified starting vector is computed following the same procedure used for the inflow case, replacing the inflow vector $I$ with the starting vector $S$.
@@ 4.4 Emission estimation model
For the calculation of emissions in the area, it is assumed that each vehicle in circulation produces a value of average emissions per kilometer depending on its characteristics. This work uses the classification of vehicles in Euro emission standards (Selleri et al., 2021), widely adopted in European environmental and transport studies, to standardize the assessment of vehicle emissions and ensure comparability across different regions and datasets. Vehicles are classified into seven Euro classes, from Euro-0 to Euro-6, based on their characteristics such as vehicle type, fuel type, date of the first registration, and maximum emissions per kilometer. Accordingly, we define $\bar{E}_{l,km}$ as the representative average emissions per kilometer corresponding to the $l$-th Euro-class, with values obtained from the literature. Moreover, let $P_{l}$ represent the percentage of vehicles for each of the Euro-class $l$. This proportion reflects the actual composition of the fleet in the current base scenario, derived from the elaboration of real-world data.
The distribution of vehicles varies before and after the implementation of the policy, as the policy imposes different costs on different Euro-classes, which in turn affects the likelihood of vehicles being rigid. Consequently, during the period of application of the policy, the Euro-level distribution of the fleet must be recalculated considering the adjusted proportions of rigid vehicles, while outside this period it must be based on the original distribution. As a first step, the rigidities for each Euro-class must be calculated. Equation 17 represents the allocation of the total rigidity fraction, $F_{rig}$, among different Euro-levels $l$. In this way, the formula distributes the total rigidity among the partitions proportionally to both the size of the partition and the likelihood of rigidity within it for each time interval $t\in[t_{s},t_{e}]$.
$F_{rig,l}(t)=F_{rig}(t)\cdot\frac{\mathbb{P}[\text{rigidity for Euro-level }l]}{\sum_{j}P_{j}\cdot\mathbb{P}[\text{rigidity for Euro-level }j]}.$ (17)
Then, it is possible to compute the Euro-level distribution under the policy enactment. During the period $t_{s},t_{e}]$, the new proportion $P_{l,m}(t)$ accounts for two contributions. First, it considers the fraction of exempt vehicles $P_{l}\cdot F_{e}$, which represents the part of the Euro-class $l$ that is not subject to the policy and remains unchanged. Second, it considers the fraction of rigid vehicles, $P_{l}\cdot\frac{F_{rig,l}(t)}{F_{e}+F_{rig}(t)}$, which represents the proportional contribution of the rigid share of class $l$ compared to the total of rigid vehicles and exempt vehicles. Equation 18 formalizes this reasoning.
@@ 5.1 Technical development
The frontend is designed to ensure accessibility and usability for domain experts, including those without technical backgrounds. To this end, we develop an interactive dashboard that allows near real-time simulation of what-if scenarios under varying conditions. Users can directly modify key control parameters, described in Sections 3 and 4, including policy fees per vehicle type, exempted vehicles, temporal windows of the policy application, behavioral parameters, and public transport settings. On the same page, simulation results are presented clearly and intuitively (see Section 5.3 for a wider discussion on qualitative and quantitative results). Global indicators summarize daily information for both the as-is baseline and what-if modified scenarios. They refer to indicators such as the total vehicle inflow per day, the maximum traffic within the day, total collected fees, and the estimated emission modification. For insight at a microscopic scale, interactive graphs are included to display the daily evolution of key quantities of interest in the current and modified scenarios. By changing the parameters, users can evaluate multiple scenarios corresponding to different combinations of parameters. The dashboard not only enables comparison between the baseline and modified scenarios, but also across different what-if scenarios (e.g., scenario 1 vs. scenario 2), supporting the assessment of alternative policy options.
@@ 5.2 Application: the case of Bologna, Italy
Figure 3: Left: the city of Bologna is highlighted in yellow, while the selected area, subject to the policy, is colored in orange. Center: population distribution in Bologna; the boundary of the selected area is highlighted in black. Right: road network and public transport stops in Bologna; the boundary of the selected area is highlighted in black.
@@
The proposed model is tested on a real-world case study in the city of Bologna in Italy, a major urban center with a large resident population and a complex multimodal mobility system. The city is already characterized by a limited traffic zone in the historical center. For experimental purposes, we simulate an additional circulation restriction applied within a broader zone: its northern boundary is delineated by the ring road and the A1-A14 motorway, while the southern boundary follows the city’s morphological constraints, namely, the hilly area. Figure 3 represents the selected area. This configuration serves as an illustrative example of how the model can be applied to large-scale area-based policies.
@@
The selected area exhibits both strong centrality and high urban attractiveness. It includes more than $321,172$ residents, 8282% of Bologna’s $392,044$ urban population [^1] Data extracted from the “Resident Population by Age, Sex, Citizenship, Neighborhood, and Statistical Area” dataset from the Comune di Bologna Open Data portal, version updated 31/12/2024; accessed 12/12/2025. URL: https://opendata.comune.bologna.it. and 3131% of the $1,051,536$ residents in the wider metropolitan province [^2] Data extracted from the “Resident Population” ISTAT dataset; accessed on 12/12/2025. URL: https://demo.istat.it.. It concentrates substantial private motorized mobility flows, with $269,226$ daily trips entering the area from outside and $329,807$ internal trips [^3] Origin-Destination data are extracted and relaborated from PUMS 2019, provided by Comune di Bologna.. The road network of the selected area covers approximately $708$ km of drivable roads, of which $287$ km correspond to primary corridors (i.e., with tags trunk, primary, secondary, and tertiary roads according to OpenStreetMap) [^4] Road network data extracted from OpenStreetMap; accessed 8/06/2025. URL: https://www.openstreetmap.org.. Public transport is also widely available in the area. In particular, for weekdays, the bus routes cover $249$ km of the drivable road network with $177$ bus lines, operating on a total of $786$ bus stops with a median frequency of $14$ trips per line per day (range: $1$ to more than 100100) [^5] Bus service data extracted from the Open Data portal of Trasporto Passeggeri Emilia‑Romagna (TPER); accessed on 25/05/2025. URL: https://solweb.tper.it/web/tools/open-data/open-data.aspx.. The area also includes Bologna Centrale, a major national and regional railway hub. The station serves high-speed and intercity trains to major Italian cities (e.g., Florence, Milan, Venice, Rome) as well as international connections, including routes to Munich, and regional and suburban lines connecting surrounding towns and the Modena-Bologna-Ferrara corridor.
@@
Bologna is further selected due to the advanced mobility data infrastructure maintained by the municipality. High-resolution data are collected from multiple sources, including traffic sensors, cameras, and vehicle registration data from the national motorization authority . These datasets allow us to estimate the current as-is traffic situation. In particular, the integration of multi-source data with statistical and data science preprocessing techniques enables the estimation of inflows into the restricted area as well as trips originating within it, with a temporal discretization of $\nu=5$ minutes. Furthermore, we have estimates of the daily percentage of vehicles by Euro emission class currently circulating in the area. From Euro-0 to Euro-6 vehicles, respectively, the proportions are $0.059$, $0.012$, $0.034$, $0.054$, $0.198$, $0.176$, and $0.467$. Similarly, knowing the vehicular fleet composition in terms of segment, fuel type, and technology, we estimate the pollutant emissions per kilometer per vehicle typology in the area. We focus on the nitrogen oxides (NOx), a group of highly reactive gases emitted by vehicles that significantly contribute to air pollution. Although NOx is our primary subject, this analytical framework can be equally applied to other components of vehicular air pollution (e.g., particulate matter [PMx], carbon monoxide [CO], carbon dioxide [CO2]). For each Euro-level class, the estimated emissions in grams per kilometer are: $0.2105847$, $0.217457$, $0.2401457$, $0.247239$, $0.135555$, $0.099559$, and $0.068246$.